# A faster FPRAS for #NFA

### 20 Dec 2023 v1

@@ 6.2. Proof of Theorem 3
[…]
thresh $=24\cdot\frac{(1+\beta)^{2\ell-2}}{\beta^{2}}\cdot\log\left(\frac{4k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}{\eta^{\prime}}\right)\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{8m^{2}n}{\delta}\right)$
$\leq\frac{384\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{8m^{2}}{\delta}\right)\leq\mathtt{ns}$
@@
${\mathsf{Pr}}\!\left[\begin{array}[]{l}\frac{|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|}{(1+\beta)^{\ell}}\leq\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\\ \quad\quad\leq(1+\beta)^{\ell}\;|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|\end{array}\big{|}\;\;\texttt{AccurateN}_{\ell-1};{{\mathsf{U}}_{\ell-1}}\right]\geq 1-\frac{\eta}{2}$
Since $\mathsf{N}({q}^{\ell})=\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$, we have
${\mathsf{Pr}}\!\left[\begin{array}[]{l}\frac{|\mathcal{L}({q}^{\ell})|}{(1+\beta)^{\ell}}\leq\mathsf{N}({q}^{\ell})\\ \quad\quad\leq(1+\beta)^{\ell}|\mathcal{L}({q}^{\ell})|\end{array}\mid\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\eta$
Lemma 6.0.
@@
${\mathsf{TV}}(S(q^{\ell}),U(q^{\ell})\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\ell-1})\leq\eta$
@@
Fix $\ell$ and $q$. Let $\mathtt{Fail}^{j}$ denote the event that the $j^{\text{th}}$ call to sample (in the loop starting at Line 4) for $i=j$ returns $\bot$. Similarly, let $\texttt{Accurate}\texttt{AppUnion}^{j}$ denote the event AccurateAppUnion corresponding to the call to sample for $i=j$. Using Iversion bracket notation, we denote the indicator variable $[\mathtt{Fail}^{j}]$. Also, we use SmallS to denote the event that $\alpha_{{q}^{\ell}}>0$. Observe that SmallS occurs if and only if $\sum\limits_{j=1}^{\mathtt{xns}}[\mathtt{Fail}^{j}]\geq\mathtt{xns}-\mathtt{ns}$.
Recall, we have, for each $j$
${\mathsf{Pr}}\!\left[[\mathtt{Fail}^{j}]\;\;\big{|}\;\;\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq 1-\frac{2}{3e^{2}}$
Therefore, from Chernoff bound, we have
${\mathsf{Pr}}\!\left[\textsc{SmallS}\;\;\big{|}\;\;\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\frac{\eta}{2}$
@@
${\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\eta/2$
Note that assuming $\texttt{AccurateN}_{\leq\ell}$ and ${\mathsf{U}}_{\leq\ell-1}$, is the event SmallS does not happen and for all $j$ $\texttt{Accurate}\texttt{AppUnion}^{j}$ holds then the distribution of $S(q^{\ell})$ and $U(q^{\ell})$ is identical. So,
@@
$\quad+{\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$
$\leq\eta/2+\eta/2=\eta$
Now, we are ready to state and prove the main theorem.
Proof of Theorem 3.
For ease of exposition, we will refer to the event $\texttt{AccurateN}_{n}$ in our analysis. Applying  Lemma 1, we have
@@
${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{n}}\right]\leq$
@@
$\sum_{\ell\in[1,n];q\in Q}{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]$
@@
$\qquad+\sum\limits_{\ell\in[1,n-1];q\in Q}\mathsf{TV}(S(q^{\ell}),U(q^{\ell})\mid\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})$
By Lemma 2 and Lemma 3 and applying union bound oven all $\ell\in[1,n]$ and all $q\in Q$ we have
@@
To calculate the time complexity of Algorithm 4 we first observe that the for loop goes over $nm$ elements. Now for each $\ell$ and $q$ there is one call to $\texttt{AppUnion}_{\beta,\eta}$ with the number of sets being $O(m)$ and $\mathtt{xns}$ calls to sample(). So by Theorem 1 the time taken by AppUnion is $O(m/\beta^{2}\log(m/\eta))$ number of calls to the oracle. So the total complexity is $\widetilde{O}(mn^{4}/\epsilon^{2}\log(1/\delta))$, ignoring the factors of $\log(n+m)$. The sample() on the other hand is a recursive function which calls $\texttt{AppUnion}_{\beta,\eta}()$ every recursion and the recursion happens $O(n)$ times. So the time complexity for each of the $\mathtt{xns}$ calls to sample() is $\widetilde{O}(mn^{5}/\epsilon^{2}\log(1/\delta))$ calls to the oracle. Now note that the membership oracle takes $O(m^{2}n)$ steps for each call. So in total the time complexity of the Algorithm 4 is $\widetilde{O}(mn\cdot\mathtt{xns}\cdot\frac{mn^{5}}{\epsilon^{2}}\log(1/\delta)\cdot(m^{2}n))=\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$.

### 08 Apr 2024 v2

@@ 4.3. Correctness of Algorithm 3
Let us denote by $\mathcal{L}(\textsf{Pred}({q}^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))$ the set $\bigcup\limits_{p\in\textsf{Pred}({q}^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})}\mathcal{L}(p).$ Observe that,
$\mathcal{L}({q}^{\ell})=\mathcal{L}(\textsf{Pred}({q}^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}))\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}\}\uplus\mathcal{L}(\textsf{Pred}({q}^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}))\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}.$
Now, we would like to apply Theorem 1 to show that $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ (at level $\ell$) is a good estimate. To this end, we first show that the pre-conditions stated in Theorem 1 hold. This is because, (1) under the event $\texttt{AccurateN}_{\leq\ell-1}$, the estimates $\{\mathsf{N}({p}^{\ell-1})\}_{p\in Q}$ satisfy desired requirements. Next, (2) for each $i\leq\ell-1$, the samples $\{\mathsf{S}({p}^{\ell-1})\}_{p\in Q}$ also obey the desired requirements. This is because thresh (in Algorithm 1), after substituting $\beta,\eta$, satisfies
thresh $=24\cdot\frac{(1+\beta)^{2\ell-2}}{\beta^{2}}\cdot\log\left(\frac{4k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}{\eta^{\prime}}\right)\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{8m^{2}n}{\delta}\right)\leq\frac{384\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{8m^{2}}{\delta}\right)\leq\mathtt{ns}$
@@
${\mathsf{Pr}}\!\left[\frac{|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|}{(1+\beta)^{\ell}}\leq\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leq(1+\beta)^{\ell}\;|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|\;\big{|}\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]$
is at least $(1-\frac{\eta}{2})$. Since $\mathsf{N}({q}^{\ell})=\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$, we have
${\mathsf{Pr}}\!\left[\frac{|\mathcal{L}({q}^{\ell})|}{(1+\beta)^{\ell}}\leq\mathsf{N}({q}^{\ell})\leq(1+\beta)^{\ell}|\mathcal{L}({q}^{\ell})|\;\big{|}\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\eta\qquad\qed$
In addition to the previously introduced notation ${\mathsf{Pr}}\!\left[;{{\mathsf{U}}_{\leq i}}\right]$, we will also use the notation ${\mathsf{TV}}(X,Y\;;{\mathsf{U}}_{\leq i})$ (resp. ${\mathsf{TV}}(X,Y\,\big{|}\,\mathcal{F}\;;{\mathsf{U}}_{\leq i})$) to denote the total variational distance between the distribution induced by random variables $X,Y$ (resp. distribution induced by $X,Y$ conditioned on the event $\mathcal{F}$) when ${\sf S}_{\leq i}$ is distribution entangled to ${\mathsf{U}}_{\leq i}$. Formally,
${\mathsf{TV}}(X,Y\mid\mathcal{F}\;;{\mathsf{U}}_{\leq i})=\sum_{\omega^{\prime}}$ ${\mathsf{Pr}}\!\left[{\mathsf{U}}_{\leq i}=\omega^{\prime}\right]\left(\sum_{\omega}\Big{(}\Pr[X=\omega\mid\mathcal{F}\cap({\sf S}_{\leq i}=\omega^{\prime})]\right.$
$-\min\left(\Pr[X=\omega\mid\mathcal{F}\cap({\sf S}_{\leq i}=\omega^{\prime})],\Pr[Y=\omega\mid\mathcal{F}\cap({\sf S}_{\leq i}=\omega^{\prime})]\right)\Big{)}\Big{)}$
In the above expression, $\omega$ ranges over the union of domains of $X$ and $Y$ while $\omega^{\prime}$ ranges over the union of domains of ${\sf S}_{\leq i}$ and ${\mathsf{U}}_{\leq i}$.
Lemma 2 ().
@@
${\mathsf{TV}}(\mathsf{S}({q}^{\ell}),\mathsf{U}({q}^{\ell}))\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})\leq\eta$
@@
Fix $\ell$ and $q$. Let $\mathtt{Fail}^{j}$ denote the event that the $j^{\text{th}}$ call to sample (in the loop starting at Line 3) for $i=j$ returns $\bot$. Similarly, let $\texttt{Accurate}\texttt{AppUnion}^{j}$ denote the event AccurateAppUnion corresponding to the call to sample for $i=j$. Using Iverson bracket notation (Iverson, 1962), we denote the indicator variable $[\mathtt{Fail}^{j}]$. Also, we use SmallS to denote the event that $\alpha_{{q}^{\ell}}>0$. Observe that SmallS occurs if and only if $\sum\limits_{j=1}^{\mathtt{xns}}[\mathtt{Fail}^{j}]\geq\mathtt{xns}-\mathtt{ns}$.
Recall, we have, for each $j$, ${\mathsf{Pr}}\!\left[[\mathtt{Fail}^{j}]\;\;\big{|}\;\;\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$ is at most $(1-\frac{2}{3e^{2}})$. Therefore, from Chernoff bound, we have
${\mathsf{Pr}}\!\left[\textsc{SmallS}\;\;\big{|}\;\;\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\frac{\eta}{2}.$
@@
${\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\frac{\eta}{2}.$
Note that given the event $\texttt{AccurateN}_{\leq\ell}$ and under the distributions obtained using ${\mathsf{U}}_{\leq\ell-1}$, if the event SmallS does not happen and the event $\bigcap_{j}\texttt{Accurate}\texttt{AppUnion}^{j}$ holds then the distribution of $\mathsf{S}(q^{\ell})$ and $\mathsf{U}(q^{\ell})$ is identical. So,
@@
$\quad\qquad\qquad+{\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\eta/2+\eta/2=\eta\qed$
4.3.1. Proof of Theorem 3
Our proof of the Theorem 3 relies on upper bounding the probability with which the estimate of Algorithm 3 is outside the desired range. For this, we will set up new random variables and relate its joint distribution with the statsitical distance between the samples constructed by the algorithm and the ideal set of uniform samples. Recall that $\{\mathsf{S}({q}^{\ell})\}_{\ell\in[0,n],q\in Q}$ represent the set of random variables corresponding to the (multi)sets of samples we obtain during run of Algorithm 3. Furthermore, $\{\mathsf{U}({q}^{\ell})\}_{\ell\in[0,n],q\in Q}$ represent the set of random variables that correspond to repeatedly sampling $\mathtt{ns}$ elements (with replacement) uniformly at random from $\mathcal{L}({q}^{\ell})$. It is worth observing that while, for all $q,p,\ell,\ell^{\prime}$, $\mathsf{S}({q}^{\ell})$ and $\mathsf{S}({p}^{\ell^{\prime}})$ are dependent but $\mathsf{U}({q}^{\ell})$ and $\mathsf{U}({q}^{\ell^{\prime}})$ are independent. Also, the definition of distribution $\mathsf{U}({q}^{\ell})$ is independent of  Algorithm 3 as it is defined solely based on the set of $\mathcal{L}(\mathsf{S}({q}^{\ell}))$. Since we are talking about the sets of random variables, there is an implicit assumption that these sets are ordered.
We will now define two sequences of random variables:
$X=\{X^{q}_{\ell}\}_{\ell\in[0,n],q\in Q}\ \mbox{ and }\ Y=\{Y^{q}_{\ell}\}_{\ell\in[0,n],q\in Q},$
Instead of explicitly specifying them, we will instead specify properties about their joint distribution. We will use the notation $X_{\leq i}$ (and resp. $Y_{\leq i})$) to denote the ordered set $\{X^{q}_{\ell}\}_{\ell\in[1,i],q\in Q}$ (resp. the ordered set $\{Y^{q}_{\ell}\}_{\ell\in[1,i],q\in Q}$). The joint distribution of $(X,Y)$ has the following properties:
  1. (1)
    For every $q\in Q$, $X^{q}_{0}=Y^{q}_{0}={\sf S}^{q}_{0}$.
  2. (2)
    For every $q\in Q,\ell\in[1,n]$, $X^{q}_{\ell}=S(q^{\ell})$.
  3. (3)
    For all sets $\omega$, sequences of sets $\omega^{\prime}$, $\ell\in[1,n]$, and $q\in Q$, we have
    ${\mathsf{Pr}}\!\left[X^{q}_{\ell}=Y^{q}_{\ell}=\omega\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]$
@@ 4.3.1. Proof of Theorem 3
    $\qquad\qquad=\min\left({\mathsf{Pr}}\!\left[U(q^{\ell})=\omega\right],{\mathsf{Pr}}\!\left[X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\right)$
    ${\mathsf{Pr}}\!\left[X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\leq{\mathsf{Pr}}\!\left[X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]$
We remark that such a joint distribution always exists. This is because (3) only restricts the case when $X_{\ell}^{q}=Y_{\ell}^{q}$, thereby, providing enough choices for assignments of probability values to the cases where $X_{\ell}^{q}\neq Y_{\ell}^{q}$ such that $Y^{q}_{\ell}$ has a well-defined probability distribution. In Appendix A.1 we give a concrete distribution that realizes these properties.
We will begin with the following simple observation, whose proof is deferred to Appendix.
Claim 3 ().
${\mathsf{Pr}}\!\left[(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega)\cap\texttt{AccurateN}_{\leq\ell-1}\right]\leq{\mathsf{Pr}}\!\left[{\mathsf{U}}_{\leq\ell-1}=\omega\right]$
Next, we relate the random variables defined above to the desired statistical distance:
Claim 4 ().
${\mathsf{Pr}}\!\left[(X_{\ell}^{q}\neq Y_{\ell}^{q})\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\cap\texttt{AccurateN}_{\leq\ell}\right]$ $\leq{\mathsf{TV}}(\mathsf{S}({q}^{\ell}),\mathsf{U}({q}^{\ell})\;\big{|}\;\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})$
Proof.
${\mathsf{Pr}}\!\left[(X_{\ell}^{q}\neq Y_{\ell}^{q})\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\cap\texttt{AccurateN}_{\leq\ell}\right]$
$=\sum_{\omega,\omega^{\prime}}{\mathsf{Pr}}\!\left[(X_{\ell}^{q}=\omega)\cap(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]$
@@
$\quad-\sum_{\omega,\omega^{\prime}}{\mathsf{Pr}}\!\left[(X_{\ell}^{q}=Y_{\ell}^{q}=\omega)\cap(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]$
$=\sum_{\omega^{\prime}}{\mathsf{Pr}}\!\left[(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\cdot$
@@
$\qquad\left(\sum_{\omega}{\mathsf{Pr}}\!\left[(X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\right.$
$\left.\qquad-\sum_{\omega}{\mathsf{Pr}}\!\left[(X_{\ell}^{q}=Y_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\right)$
$\leq\sum_{\omega^{\prime}}{\mathsf{Pr}}\!\left[U_{\leq\ell-1}=\omega^{\prime}\right]\cdot\big{(}\sum_{\omega}\left({\mathsf{Pr}}\!\left[(X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\right.$
$\qquad\qquad-\min\left({\mathsf{Pr}}\!\left[U(q^{\ell})=\omega\right],\left.\left.{\mathsf{Pr}}\!\left[X_{\ell}^{q}=\omega\mid(X_{\leq\ell-1}=\omega^{\prime})\cap\texttt{AccurateN}_{\leq\ell}\right]\right)\right)\right)$
$={\mathsf{TV}}(\mathsf{S}({q}^{\ell}),\mathsf{U}({q}^{\ell})\;\big{|}\;\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})$
We now focus on deriving an upper bound on $\texttt{AccurateN}_{n}$ as stated in the following claim, whose proof is deferred to Appendix.
Claim 5 ().
$\overline{\texttt{AccurateN}_{n}}\subseteq\bigcup_{\ell=1}^{n-1}(\overline{\texttt{AccurateN}_{\ell}}\cap{\texttt{AccurateN}_{\leq\ell-1}}\cap(X_{\leq\ell-1}=Y_{\leq\ell-1}))$
$\qquad\qquad\qquad\qquad\cup\bigcup_{\ell=1}^{n-1}\left((X_{\ell}\neq Y_{\ell})\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\cap{\texttt{AccurateN}_{\leq\ell}}\right)$
We can now finish the proof of Theorem 3. We first observe that:
${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{\ell}}\cap{\texttt{AccurateN}_{\leq\ell-1}}\cap(X_{\leq\ell-1}=Y_{\leq\ell-1})\right]$
$\leq\sum_{\omega,q}{\mathsf{Pr}}\!\left[\texttt{AccurateN}_{\leq\ell-1}\cap(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega)\right]\times$
$\qquad\qquad\qquad\qquad{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega)\cap{\texttt{AccurateN}_{\leq\ell-1}}\right]$
$\leq\sum_{\omega,q}{\mathsf{Pr}}\!\left[U_{\leq\ell-1}=\omega\right]{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\mid(X_{\leq\ell-1}=Y_{\leq\ell-1}=\omega)\cap{\texttt{AccurateN}_{\leq\ell-1}}\right]$
$=\sum_{\omega,q}{\mathsf{Pr}}\!\left[\texttt{AccurateN}_{\leq\ell-1}=\omega\right]{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\mid({\sf S}_{\leq\ell-1}=\omega\cap Y_{\leq\ell-1}=\omega\cap{\texttt{AccurateN}_{\leq\ell-1}}\right]$
$\leq\sum_{q}{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]$
The final step in the derivation uses the observation that given ${\sf S}_{\leq\ell-1}$ and ${\texttt{AccurateN}_{\leq\ell-1}}$, the probability of the event $\texttt{AccurateN}_{q,\ell}$ is only dependent on the randomness of the algorithm AppUnion. Thus, given ${\sf S}_{\leq\ell-1}$ and ${\texttt{AccurateN}_{\leq\ell-1}}$, the event $\texttt{AccurateN}_{q,\ell}$ is conditionally independent of $Y_{\leq\ell-1}$.
Putting together, we have:
${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{n}}\right]\leq$ $\sum_{\ell\in[1,n];q\in Q}{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]$
$\qquad+\sum\limits_{\ell\in[1,n-1];q\in Q}{\mathsf{TV}}(\mathsf{S}({q}^{\ell}),\mathsf{U}({q}^{\ell})\mid\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})$
By Lemma 1 and Lemma 2 and applying union bound over all $\ell\in[1,n]$ and all $q\in Q$ we have
@@
To calculate the time complexity of Algorithm 3 we first observe that the for loop goes over $nm$ elements. Now for each $\ell$ and $q$ there is one call to $\texttt{AppUnion}_{\beta,\eta}$ with the number of sets being $O(m)$ and $\mathtt{xns}$ calls to sample(). So by Theorem 1, the time taken by AppUnion is the time it takes to do $O(m/\beta^{2}\log(m/\eta))$ number of calls to the membership oracle. So the number of calls to the membership oracle is $\widetilde{O}(mn^{4}/\epsilon^{2}\log(1/\delta))$, ignoring the factors of $\log(n+m)$. The method sample() on the other hand is a recursive function which calls $\texttt{AppUnion}_{\beta,\eta}()$ in every recursive step and the depth of the recursion is $O(n)$. So the time complexity for each of the $\mathtt{xns}$ calls to sample() is the time it takes for $\widetilde{O}(mn^{5}/\epsilon^{2}\log(1/\delta))$ calls to the oracle. Now, we note that the time complexity of the membership calls can be amortized as follows. First, for every string $w\in\bigcup\limits_{q,\ell}\{\mathsf{S}({q}^{\ell})\}$, we pre-compute and store the set of reachable states of $w$ in $\mathcal{A}$ in time $m^{2}|w|\in O(m^{2}n)$ time. This takes a total of $\widetilde{O}(m^{2}n\cdot mn^{5}/\epsilon^{2}\log(1/\delta))=\widetilde{O}(m^{3}n^{6}/\epsilon^{2}\log(1/\delta))$ time. The membership checks (in Algorithm 1, Line 1) can now be performed in $O(1)$ time. So in total the time complexity of the Algorithm 3 is $\widetilde{O}(mn\cdot\mathtt{xns}\cdot\frac{mn^{5}}{\epsilon^{2}}\log(1/\delta)+m^{3}n^{6}/\epsilon^{4}\log(1/\delta))$ which is $\widetilde{O}((m^{2}n^{10}+m^{3}n^{6})\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$.